I'm in need of help very badly I have read and reread the assignment looked at the example.And I still don't understand I talked to my professor done what she said to do and nope it's a lost cause i'm starting to think , I have used tutors
but we are only allowed so many hours. here's the question.
6
--
7
+
3 3
--
7
------

Answer by Alan3354(69443)

Do you mean 6/7 + 3 3/7 ?
The denominators are the same, you can add
= 3 3/7 is 3 + 3/7
-----
6/7 + 3 + 3/7 = 3 + 9/7
9/7 = 7/7 + 2/7 = 1 + 2/7
---
3 + 1 + 2/7 = 4 + 2/7
or 4 2/7
--------------
Enter fractions as I did, not as you did.